# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v4

@@ IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$
When the state $\xi$ enters in the jump set $\mathcal{J}_{0}$ (20), the update law $\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$ is given as
@@
where $\tau_{s}>0$. Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, the set-valued mapping $\mathbf{A}:\mathbb{R}^{3}\rightrightarrows\mathbb{S}^{2}$ is defined as
$\mathbf{A}(\mathbf{x})=\begin{cases}\mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}=\mathbf{0},\\ \mathbf{q}\in\pm\frac{\mathbf{x}^{\times}\mathbf{x}_{\pi}}{\lVert\mathbf{x}^{\times}\mathbf{x}_{\pi}\rVert},&\mathbf{x}^{\times}\mathbf{x}_{\pi}\neq\mathbf{0},\end{cases}$ (28)
where for any $\mathbf{p}\in\mathbb{R}^{3}$, the set $\mathcal{P}^{\perp}(\mathbf{p})$, which is defined as
$\mathcal{P}^{\perp}(\mathbf{p}):=\{\mathbf{q}\in\mathbb{S}^{2}|\mathbf{q}^{\intercal}\mathbf{p}=0\},$ (29)
@@
According to (27), when the robot switches from the move-to-target mode to the obstacle-avoidance mode, the coordinates of the hit point $\mathbf{h}$ get updated to the current value of $\mathbf{x}.$ Moreover, according to (28), the unit vector $\mathbf{a}$ is updated such that it is perpendicular to the vector $\mathbf{h}$ and the resulting hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$ intersects with the interior of the nearest $r_{a}-$dilated obstacle. Since the target location $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a})$ and the hit point $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}$, the hyperplane intersects with both the landing region $\mathcal{R}_{l}$ and the exit region $\mathcal{R}_{e}$ associated with the nearest obstacle. This feature is then used to ensure that the robot, operating in the obstacle-avoidance mode, always enters in the move-to-target mode. This allows one to ensure global convergence to the target location, as stated later in Theorem 1.
When the state $\xi$ enters in the jump set $\mathcal{J}_{1}$, defined in (24), the update law $\mathbf{L}_{1}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)$ is given by
@@
Since the parameter $\gamma\in(r_{s},\delta-r_{s})$, according to Assumption 8, the $(r_{a}+\gamma)-$dilated obstacles $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, $\forall i\in\mathbb{I}$, are disjoint, and the target location $\mathbf{0}\in\mathcal{W}_{r_{a}+\gamma}^{\circ}$. Furthermore, according to (23), the set $\mathcal{F}_{1}^{\mathcal{W}}$ is contained within the region $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$. Hence, the proposed control law enables the robot to avoid one obstacle at a time.
@@ V STABILITY ANALYSIS
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}(\xi),\;{\xi\in\mathcal{F}},}\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}(\xi),\;{\xi\in\mathcal{J}},}$ (31)
@@
The next lemma shows that the hybrid closed-loop system (31) satisfies the hybrid basic conditions [20, Assumption 6.5], which guarantees the well-posedness of the hybrid closed-loop system.
@@
    the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, defined in (25), are closed subsets of $\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}\times\mathbb{R}$.
@@
    the flow map $\mathbf{F}$, defined in (31), is outer semicontinuous and locally bounded relative to $\mathcal{F}$, dom $\mathcal{F}\subset\text{dom }\mathbf{F},$ and $\mathbf{F}(\xi)$ is convex for every $\xi\in\mathcal{F}$.
@@
Under Assumption 8, for the hybrid closed-loop system (31), the obstacle-free set $\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{2}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ is forward invariant.
@@
Next, we provide one of our main results which establishes the fact that for all initial conditions in the obstacle-free set $\mathcal{K}$, the proposed hybrid controller not only ensures safe navigation but also guarantees global asymptotic stability of the target location at the origin.

### v5

@@ IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$
When the state $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode, defined in (26), the update law $\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$ is given as
@@
where $\tau_{s}>0$. Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, the set-valued mapping $\mathbf{A}$ is defined as
$\mathbf{A}(\mathbf{x})=\begin{cases}\mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}=\mathbf{0},\\ \mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x})\cap\mathcal{P}(\mathbf{x},\mathbf{x}_{\pi}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}\neq\mathbf{0},\end{cases}$ (35)
where for any $\mathbf{p}\in\mathbb{R}^{n}$, the set $\mathcal{P}^{\perp}(\mathbf{p})$, which is defined as
$\mathcal{P}^{\perp}(\mathbf{p}):=\{\mathbf{q}\in\mathbb{S}^{n-1}|\mathbf{q}^{\intercal}\mathbf{p}=0\},$ (36)
@@
As per (34), when $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode, the current value of $\mathbf{x}$ is assigned to the hit point $\mathbf{h}$. Moreover, using (35), the unit vector $\mathbf{a}$ is updated to be perpendicular to $\mathbf{x}$ and lies in the plane spanned by $\mathbf{x}$ and $\mathbf{x}_{\pi}$, provided they are not collinear. As a result, the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ passes through the target location at the origin and intersects the interior of the obstacle being avoided. This ensures that $\mathcal{P}(\mathbf{h},\mathbf{a})$ intersects both the avoidance region $\mathcal{R}_{a}$ and the exit region $\mathcal{R}_{e}$ associated with the obstacle to be avoided. This property guarantees that, while operating in the obstacle-avoidance mode, the robot eventually enters in the exit region where the obstacle no longer blocks its straight path to the origin, and switches back to the move-to-target mode. This allows one to establish global convergence properties of the target location at the origin, as stated later in Theorem 1.
When the state $\xi$ enters in the jump set $\mathcal{J}_{1}$ of the obstacle-avoidance mode, defined in (31), the update law $\mathbf{L}_{1}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)$ is given by
@@
Since the parameter $\gamma\in(0,\delta-r_{s})$, according to Assumption 9, the $(r_{a}+\gamma)-$dilated obstacles $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, $\forall i\in\mathbb{I}$, are disjoint. Furthermore, according to (30), the set $\mathcal{F}_{1}^{\mathcal{W}}$ is contained within the region $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$. Hence, the proposed control law enables the robot to avoid one obstacle at a time.
Control design summary: The proposed hybrid feedback control law can be summarized as follows:
  • Parameters selection: the target location is set at the origin with $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ}$, and $\xi(0,0)\in\mathcal{K}$. The gain parameters $\kappa_{s}$ and $\kappa_{r}$ are set to positive values, and a sufficiently small positive value is chosen for $\bar{\epsilon}$, used in (28). The scalar parameter $\gamma$, used in the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, is selected such that $\gamma\in(0,\delta-r_{s})$. The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$ The parameter $\tau_{s}$ is selected such that $\tau_{s}>0$ and $\delta_{s}$, used in (31), is set to ensure that $0<\delta_{s}<\tau_{s}$.
  • Move-to-target mode $m=0$: this mode is activated when $\xi\in\mathcal{F}_{0}$. As per (12a), the control input is given by $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x}$, causing $\mathbf{x}$ to evolve along the line segment $\mathcal{L}_{s}(\mathbf{0},\mathbf{x})$ towards the origin. If, at some instance of time, $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode, the state variables $(\mathbf{h},\mathbf{a},m,s)$ are updated using (34), and the control input switches to the obstacle-avoidance mode.
  • Obstacle-avoidance mode $m=1$: this mode is activated when $\xi\in\mathcal{F}_{1}$. As per (12a), the control input is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, causing $\mathbf{x}$ to evolve in the $\gamma-$neighborhood of the nearest obstacle along the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ until the state $\xi$ enters in the jump set $\mathcal{J}_{1}$ of the obstacle-avoidance mode. When $\xi\in\mathcal{J}_{1}$, the state variables $(m,s)$ are updated using (37), and the control input switches to the move-to-target mode.
@@ V Stability Analysis
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}(\xi),{\xi\in\mathcal{F}}}\quad\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}(\xi),{\xi\in\mathcal{J}}},$ (38)
@@
The next lemma shows that the hybrid closed-loop system (38) satisfies the hybrid basic conditions [27, Assumption 6.5], which guarantees the well-posedness of the hybrid closed-loop system.
@@
    the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, defined in (32), are closed subsets of $\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}$.
@@
    the flow map $\mathbf{F}$, defined in (38), is outer semicontinuous and locally bounded relative to $\mathcal{F}$, dom $\mathcal{F}\subset\text{dom }\mathbf{F},$ and $\mathbf{F}(\xi)$ is convex for every $\xi\in\mathcal{F},$
@@
Under Assumption 9, for the hybrid closed-loop system (38), the obstacle-free set $\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ is forward invariant.
@@
Next, we provide one of our main results which states that, for all initial conditions in the obstacle-free set $\mathcal{K}$, the proposed hybrid controller not only ensures safe navigation but also guarantees global asymptotic stability of the target location at the origin.
Theorem 1.
  1. i)
    the obstacle-free set $\mathcal{K}$ is forward invariant;
  2. ii)
    the target set $\mathcal{A}$ is globally asymptotically stable over the set $\mathcal{K}$;
  3. iii)
    the number of jumps is finite.
Proof.
See Appendix -E. ∎
VI Application to sphere worlds
Obviously, the hybrid feedback controller (12), which is designed for safe autonomous navigation in $n-$D environments with arbitrary convex obstacles, is applicable in sphere words i.e., environments with spherical obstacles. However, in this section, we will take advantage of the simplified geometry of the obstacles to redesign the control law (12) in a way that ensures a monotonic decrease of the distance between the robot and the target location–a feature that is not guaranteed with the control law (12) in environments with arbitrary convex obstacles.
Let us consider environments with spherical obstacles as defined in [17]. The workspace $\mathcal{W}:=\mathcal{B}_{r_{0}}(\mathbf{c}_{0})$ is a compact sphere with radius $r_{0}\in\mathbb{R}_{>0}$ and center $\mathbf{c}_{0}\in\mathbb{R}^{n}.$ In addition, the workspace $\mathcal{W}$ contains disjoint, compact spherical obstacles $\mathcal{O}_{i}:=\mathcal{B}_{r_{i}}(\mathbf{c}_{i}),i\in\mathbb{I}\setminus\{0\}$, where $r_{i}\in\mathbb{R}_{\geq 0}$ and $\mathbf{c}_{i}\in\mathcal{W}$ represent the radius and the center of obstacle $\mathcal{O}_{i}.$ Similar to [17], the workspace $\mathcal{W}$ satisfies Assumption 9.
Taking advantage of the spherical geometry of the obstacles, we will appropriately design the unit vector $\mathbf{a}$ and modify the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ in (13), to ensure that in the obstacle-avoidance mode, the distance between the target location and the robot is monotonically decreasing.
Figure 7: The left figure depicts the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{P}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, shaded in a dark blue color. The right figure shows the direction of the modified obstacle-avoidance control vector, $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, at $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{P}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$.
For any $\mathbf{p}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, we define a set-valued mapping $\mathbf{A}_{s}$ as follows:
$\mathbf{A}_{s}(\mathbf{p})=\{\mathbf{q}\in\mathbf{A}(\mathbf{p})|\mathbf{q}^{\top}\mathbf{p}_{\pi}\geq 0\},$ (40)
where the set-valued mapping $\mathbf{A}$ is defined in (35), and $\mathbf{p}_{\pi}=\mathbf{p}-\Pi(\mathbf{p},\mathcal{O}_{\mathcal{W}})$.
Note that for any hit point $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for $i\in\mathbb{I}$, if one sets $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$, then, since obstacle $\mathcal{O}_{i}$ is a sphere, it can be shown that for all $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ $\Pi(\mathbf{x},\mathcal{O}_{i})$ belongs to the plane $\mathcal{P}(\mathbf{h},\mathbf{a}).$ Therefore, for any given $\mathbf{h}\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})$, $\beta\in[r_{a},r_{a}+\gamma]$ and for all $\mathbf{x}\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one can show that the vector $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ belongs to the intersection of the hyperplane $\mathcal{H}(\mathbf{0},\mathbf{x}_{\pi})$ and the plane $\mathcal{P}(\mathbf{h},\mathbf{a}),$ where $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$. This allows one to ensure that if $\mathbf{h}\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})$ for $\beta\in[r_{a},r_{a}+\gamma]$, and $\mathbf{x}(t_{0},j_{0})\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for some $(t_{0},j_{0})\in\text{ dom }\xi$, where $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$, then under the control input $\mathbf{u}(\xi)=\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, where
$\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})=\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},$ (41)
one has $\mathbf{x}(t,j)\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all time $(t,j)\succeq(t_{0},j_{0}),$ ensuring robot’s safety, as stated later in Theorem 2
Now, for a given $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, consider the hyperplane $\mathcal{H}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, where $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$ and $\mathbf{c}_{i}$ is the center of obstacle $\mathcal{O}_{i}$. Notice that, for all locations $\mathbf{x}$ in the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, which is depicted in Fig. 7, the inner product between the modified obstacle-avoidance control term $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$ and the vector $\mathbf{x}$ is always non-positive. This allows one to show that if $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$ and $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, then the control input vector $\mathbf{u}(\xi)=\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$ will ensure monotonic decrease of the distance $\|{\mathbf{x}}\|$ as long as the state $\mathbf{x}$ remains in the set $\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, as stated later in Theorem 2.
Next, motivated by the preceding discussion, we modify the proposed hybrid feedback control law (12) using the modified obstacle-avoidance control vector $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, defined in (41), and the set-valued mapping $\mathbf{A}_{s}$, provided in (40), as follows:
  1. 1.
    the modified hybrid control input vector $\mathbf{u}_{s}(\xi)$, which is obtained by replacing the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ with the modified obstacle-avoidance control vector $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$ in (12a), is given as
    $\mathbf{u}_{s}(\xi)=-\kappa_{s}(1-m)\mathbf{x}+\kappa_{r}m\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a}),$ (42)
    where $\kappa_{s}>0$ and $\kappa_{r}>0.$
  2. 2.
    the modified update law $\mathbf{L}_{0}^{s}(\xi)$ when the state $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode is given as
    $\mathbf{L}_{0}^{s}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=\left\{\begin{bmatrix}\mathbf{x}\\ \mathbf{a}^{\prime}\\ 1\\ s+\tau_{s}\end{bmatrix},\mathbf{a}^{\prime}\in\mathbf{A}_{s}(\mathbf{x})\right\},$ (43)
    where the set-valued mapping $\mathbf{A}_{s}$ is defined according to (40).
The hybrid closed-loop system resulting from the modified hybrid feedback control law is given by
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}_{s}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}_{s}(\xi),{\xi\in\mathcal{F}}}\quad\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}_{s}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}_{s}(\xi),{\xi\in\mathcal{J}}},$ (44)
where the control input vector $\mathbf{u}_{s}$ is defined in (42) and the update law $\mathbf{L}_{s}$ is obtained by replacing $\mathbf{L}_{0}$ in (33) with $\mathbf{L}_{0}^{s}$ given in (43).
The next lemma shows that the hybrid closed-loop system (44) with the data $(\mathcal{F},\mathbf{F}_{s},\mathcal{J},\mathbf{J}_{s})$ satisfies the hybrid basic conditions, as stated in Lemma 3.
Lemma 5.
The hybrid closed-loop system (44) with data $(\mathcal{F},\mathbf{F}_{s},\mathcal{J},\mathbf{J}_{s})$ satisfies the hybrid basic conditions stated in Lemma 3.
The proof of Lemma 5 is similar to the proof of Lemma 3, therefore, it is omitted.
Next, we demonstrate that for the robot operating in a sphere world, which satisfies Assumption 9, the modified proposed hybrid feedback controller ensures safe navigation. It also guarantees global asymptotic stability of the target location at the origin, with a monotonic decrease in the distance between the robot’s center and the target location.